# Noether’s theorem for the conditional principle of least action

### v1

@@
Abstract.
We propose an extension of the first Noether’s theorem to the theories whose field equations result from the requirement of conditional extremum of the action. The key ingredient of the extension is the gauge symmetry of the differential equations constraining the admissible class of field configurations. We consider a special type of global symmetries of the action which we call conditional symmetries. Such global symmetries must be special cases of gauge transformations of the constraint equations. We construct conservation laws that follow from the conditional symmetries of action. We also prove the converse theorem which connects the conserved currents to the conditional symmetries of action. The general method is illustrated by several mechanical and field-theoretical examples.
1. Introduction
We consider the problem of the conditional principle of least action, where the class of admissible field configurations is restricted by differential equations. It is known that the inclusion of constraints with Lagrange multipliers provides the necessary conditions for the conditional extremum, but the multipliers can introduce additional degrees of freedom that are not related to the original dynamics [^1] For holonomic constraints, Lagrange multipliers do not introduce any additional degree of freedom and can be algebraically eliminated from the equations. There is another notable exception: equations in which all degrees of freedom are gauged out when the equations are considered without any connection to the action. Including these equations with Lagrange multipliers in the action does not introduce new degrees of freedom. In the case of general differential equations, Lagrange multipliers create new degrees of freedom.. The conserved Noether currents for the action with Lagrange multipliers will involve these non-physical degrees of freedom. So, these conserved currents have no direct physical interpretation from the perspective of the original conditional extremum problem. We propose another idea for constructing conserved quantities related to global symmetries of the conditional extremum problem, so that the resulting currents involve only the original degrees of freedom.
@@ 1. Introduction
Given the vast body of literature on Noether’s theorems, a general overview of this topic is beyond the scope of this article, which deals specifically with the problem of conditional extremum. For the review of the general area of Noether’s theorem and detailed bibliography we refer to the monograph [2]. We can also note that if the field equations do not follow from the unconditional variational principle but admit some additional structures, then a one-way relationship between global symmetries and conservation laws can be possible. For example, a presymplectic structure [3], [4] maps symmetries to conserved currents, [5], [6], [7], [8], and a Lagrange anchor [9] defines an inverse mapping from conservation laws to symmetries [10]. In this paper, we construct a generalization of the first Noether theorem, and the converse one, for the case of the conditional extremum problem. We proceed from the usual definitions and assumptions related to the issues of symmetries and conserved currents, known from various textbooks on field theory, see, for example, [11], while for the gauge symmetry we use the standard setup and notation of gauge theories [12].
The article is organised as follows. In the next section, we briefly outline the conditional Lagrange equations that follow from the conditional principle of least action. The conditional extremum of the action is sought in the class of admissible fields that obey differential equations. This Section also addresses the gauge symmetry of the equations that constrain the fields. In Section 3, we consider special infinitesimal global symmetries of the action in the class of the fields subject to differential equations. These conditional symmetries of action represent specializations of gauge transformations of the equations that are imposed to constrain the admissible class of fields. We will then show that such conditional global symmetries lead to currents that are conserved at conditional extremals of the action. This fact is a main result of the article. We also prove the converse theorem: Every current, which is conserved at conditional extremals of the action, leads to a conditional symmetry.Throughout this paper, we develop the formalism in its full generality, without restricting the order of derivatives. At the end of Section 3, we provide a brief exposition of the key relations, including an explicit construction of conserved currents for the simplest case where the Lagrangian does not include higher-order derivatives of the fields, and gauge transformations of the constraint equations do not involve higher-order derivatives of the gauge parameters. In Section 4, we illustrate the general formalism of Sections 2-3 by examples. The final section discusses the results and possible areas of application of the proposed formalism, and more broadly, of conditionally Lagrangian theories in general.
2. Field equations for the problem of conditional extremum
@@
(1) $S[\phi(x)]=\int d^{d}xL(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)$
and impose the differential equations
(2) $T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0,\quad a=1,\ldots,N$
to restrict the admissible fields. The problem is to find the minimum of the action (1) among the fields that are solutions of equations (2).
@@
(3) $S_{\lambda}[\phi(x),\lambda(x)]=\int dx\left(L(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)+\lambda^{a}T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)\right)\,.$
@@
(4) $\frac{\delta S}{\delta\phi^{i}}=0\,,\quad\frac{\delta S}{\delta\lambda^{a}}\equiv T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0\,$
@@
(6) $\sum_{k=0}\left(\frac{\partial T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)}{\partial\,\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\phi^{i}}\right)_{\big|T=0}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\delta^{\prime}\phi^{i}=0\,.$
@@
(7) $\delta^{\prime}S=\int dx\left(\frac{\delta S}{\delta\phi^{i}}\delta^{\prime}\phi^{i}+total\,\,divergence\,\right)=0,$
and for field configuration admitted by the constraints (2). The above relation does not immediately lead to a system of differential equations, unlike the variational principle for the action with Lagrange multipliers (4), because the variations $\delta^{\prime}\phi^{i}$, being restricted by equations (6), are not arbitrary functions. As a result, the infinitesimal symmetries of the action do not necessarily lead to conserved currents even if equations (2) share the same symmetry with $S$.
Now let us explain, following [1], how to explicitly construct the field equations that are necessary conditions for the action extremum in the class of fields constrained by differential equations. First, we should find the infinitesimal gauge symmetry transformations for equations (2) which are imposed to select the admissible field configurations. It is assumed that the transformation includes derivatives of the gauge parameters $\epsilon^{\alpha}(x),\alpha=1,\ldots,m$, which are arbitrary smooth functions, up to a certain finite order $K$
(8) $\delta_{\epsilon}\phi^{i}=\hat{R}^{i}_{\alpha}\epsilon^{\alpha}=\sum_{k=0}^{K}\stackrel{{\scriptstyle(k)}}{{R}}{}^{i}_{\alpha}{}^{\mu_{1}\ldots\mu_{k}}(x,\phi,\partial\phi,\partial^{2}\phi\ldots)\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\epsilon^{\alpha}\,$,
where $\stackrel{{\scriptstyle(k)}}{{R}}{}^{i\mu_{1}\ldots\mu_{k}}_{\alpha}$ are smooth functions of the fields and their derivatives. The above transformation is supposed to be a symmetry of equations (2), hence
(9) $\delta_{\epsilon}T_{a}{}_{\big|T=0}=0,\quad\forall\epsilon_{\alpha}(x)\,.$
@@
(11) $\delta S[\phi(x)]=\int dx\left(\partial_{\mu}J^{\mu}+\frac{\delta S}{\delta\phi^{i}}\delta\phi^{i}\right)\,$,
(12) $J^{\mu}=\sum_{p=1}^{M}p^{\mu\mu_{1}\ldots\mu_{p-1}}_{i}\partial_{\mu_{1}}\ldots\partial_{\mu_{p-1}}\delta\phi^{i}\,$,
(13) $p^{\mu\mu_{1}\ldots\mu_{p-1}}_{i}=\sum_{k=0}^{M-p}(-1)^{k}\partial_{\nu_{1}}\ldots\partial_{\nu_{k}}\frac{\partial L}{\partial(\partial_{\nu_{1}}\ldots\partial_{\nu_{k}}\partial_{\mu}\partial_{\mu_{1}}\partial_{\mu_{p-1}}\phi^{i})}\,$,
@@
(14) $\delta_{\epsilon}S=\int dx\left(\epsilon^{\alpha}\hat{R}^{\dagger i}_{\alpha}\,\frac{\delta S}{\delta\phi^{i}}+\partial_{\mu}I^{\mu}\right)\,$,
@@
(15) $\hat{R}^{\dagger i}_{\alpha}F=\sum_{k=0}^{K}(-1)^{k}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\left(\stackrel{{\scriptstyle(k)}}{{R}}{}^{i\mu_{1}\ldots\mu_{k}}_{\alpha}F\right),$
@@
(17) $\frac{\delta_{\epsilon}S}{\delta\epsilon^{\alpha}}=\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}=0\,,\qquad T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0\,.$
@@
3. Nother’s theorem for the problem of conditional extremum
@@
(18) $\epsilon^{\alpha}_{\omega}(x)=\psi^{\alpha}_{A}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)\,\omega^{A}\,,\quad A=1,\dots,s\,$,
@@
(19) $\tilde{\delta}_{\omega}\phi^{i}=\delta_{\epsilon}\phi^{i}_{|\epsilon\,\mapsto\,\epsilon_{\omega}}$
@@
(20) $\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}\Lambda_{A}^{\mu}(\phi,\partial\phi,\dots)+\hat{W}^{a}_{A}T_{a}\right),$
@@
(21) $\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}I^{\mu}_{A}+\psi^{\alpha}_{A}\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}\right)\,$,
@@
(23) $\partial_{\mu}\tilde{I}{}^{\mu}_{A}=-\psi^{\alpha}_{A}\hat{R}{}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}+\hat{W}^{a}_{A}T_{a}\,,\quad\tilde{I}{}^{\mu}_{A}=I^{\mu}_{A}-\Lambda^{\mu}_{A}\,.$
@@
(24) $\partial_{\mu}{\Phi}{}^{\mu}_{A}=\hat{\Psi}{}^{\alpha}_{A}\hat{R}{}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}+\hat{W}^{a}_{A}T_{a}\,$,
@@
(25) $\hat{\Psi}^{\alpha}_{A}=\sum_{k=0}^{N}\stackrel{{\scriptstyle(k)}}{{\psi}}{}^{\alpha\,\mu_{1}\ldots\mu_{k}}_{A}(\phi,\partial\phi,\partial^{2}\phi\ldots)\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\,,\qquad\hat{\Psi}{}^{\alpha}_{A}{}_{\big|T=0}\neq 0,\quad\forall A\,.$
@@
(26) $\partial_{\mu}\Phi^{\prime\mu}_{A}=\hat{W}^{a}_{A}T_{a}+\psi^{\alpha}_{A}\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}},$
@@
(27) $),\Phi^{\prime\mu}_{A}=\Phi^{\mu}_{A}-\sum_{p=0}^{N-1}\bigg(\partial_{\mu_{1}}\ldots\partial_{\mu_{p}}\left(\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}\right)\sum_{k=0}^{N-p-1}\bigg(\partial_{\mu_{p+1}}\ldots\partial_{\mu_{p+k}}\left((-1)^{k}\stackrel{{\scriptstyle(k+p+1)}}{{\psi}}{}^{\alpha}_{A}{}^{\mu\mu_{1}\ldots\mu_{p}\mu_{p+1}\ldots\mu_{p+k}}\right)\bigg$$\bigg)\,$,
(28) $\psi^{\alpha}_{A}=\sum_{k=0}^{N}(-1)^{k}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\stackrel{{\scriptstyle(k)}}{{\psi}}{}^{\alpha}_{A}{}^{\mu_{1}\ldots\mu_{k}}\,$,
@@
(29) $\delta_{\epsilon}\phi^{i}=\hat{R}^{i}_{\alpha}\epsilon^{\alpha},\quad\epsilon_{\omega}=\psi^{\alpha}_{A}\omega^{A}\,,\quad\tilde{\delta}_{\omega}\phi^{i}=\delta_{\epsilon_{\omega}}\phi^{i}\,.$
@@
(30) $\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}\left(I^{\mu}_{A}+\Phi^{\prime\mu}_{A}\right)+\hat{W}{}^{a}_{A}T_{a}\right)\,.$
@@
The formalism of the previous two sections covers the general case of Lagrangian, which involves the derivatives of the fields of any order, and the gauge transformations (8) with any derivatives of gauge parameters. The full generality of the formalism may leave the impression that the construction of conditionally Lagrangian equations and the conserved currents is a rather cumbersome procedure. However, Lagrangians without higher-order derivatives are most often studied, and gauge symmetries also most often include only the first derivatives of the parameters. In such a situation, without higher-order derivatives, both the conditional Lagrangian equations and the corresponding conditional symmetries and conserved currents take on a simpler form. Therefore, for the convenience of those readers who might be more interested in searching for conditional conservation laws in first-order models than in the generality of the formalism, we present a summary of the key expressions for this simpler type of theory.
Let us assume that gauge transformation of the constraining equations $T(\phi,\partial\phi,\partial^{2}\phi,\dots)=0$ involves at most the first derivatives of the gauge parameters,
(31) $\delta_{\epsilon}\phi^{i}=\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}(\phi,\partial\phi,\partial^{2}\phi,\dots)\epsilon^{\alpha}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}(\phi,\partial\phi,\partial^{2}\phi,\dots)\partial_{\mu}\epsilon^{\alpha}\,,\quad\delta_{\epsilon}T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)_{\big|T=0}=0\,.$
Let us further assume that the Lagrangian does not involve higher derivatives. Then the conditional extremals of action are subject to the following conditionally Lagrangian equations,
(32) $\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\mu}\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\right)-\partial_{\mu}\left(\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\mu}\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\right)\right)=0\,,\quad T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)=0\,.$
@@
(33) $\tilde{\delta}_{\omega}\phi^{i}=\left(\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\psi^{\alpha}_{A}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\partial_{\mu}\psi^{\alpha}_{A}\right)\omega^{A}\,.$
@@
(34) $\tilde{\delta}_{\omega}L_{\big|T=0}=\omega^{A}\partial_{\mu}\Lambda^{\mu}_{A}(\phi,\partial\phi,\dots)\,.$

### v2

@@
Abstract
We consider the problem of a conditional extremum of an action in a class of fields constrained by differential equations. For this setup, we propose an extension of Noether’s first theorem to connect the symmetries of the action and the imposed equations to the currents conserved at the conditional extrema. The key ingredient of the extension is the gauge symmetry of the differential equations constraining the admissible class of field configurations. We consider a special type of global symmetries of the action which we call conditional symmetries. Such global symmetries must be special cases of gauge transformations of the constraint equations. We construct conservation laws that follow from the conditional symmetries of action. No Lagrange multipliers or other auxiliary fields are introduced and the conserved currents include only the original fields. We also prove the converse theorem which connects the conserved currents to the conditional symmetries of action. The general method is illustrated by several mechanical and field-theoretical examples.
1 Introduction
We consider the problem of the conditional principle of least action, where the class of admissible field configurations is restricted by differential equations. It is known that the inclusion of constraints with Lagrange multipliers provides the necessary conditions for the conditional extremum of the functional, but the multipliers can introduce additional degrees of freedom that are not related to the original dynamics [^1] For holonomic constraints, Lagrange multipliers do not introduce any additional degree of freedom and can be algebraically eliminated from the equations. Another notable exception: equations in which all degrees of freedom are gauged out when the constraints are considered without any connection to the action. Including these equations with Lagrange multipliers in the action does not introduce new degrees of freedom. For general differential equations, Lagrange multipliers create new degrees of freedom.. The conserved Noether currents for the action with Lagrange multipliers will involve these non-physical degrees of freedom. So, these conserved currents have no direct interpretation from the perspective of the original conditional extremum problem. We propose a different idea for constructing conserved quantities related to global symmetries of the conditional extremum problem, so that the resulting currents involve only the original degrees of freedom.
@@ 1 Introduction
Given the vast body of literature on Noether’s theorems, a general overview of this topic is beyond the scope of this article, which deals specifically with the problem of conditional extremum. For the review of the general area of Noether’s theorem and detailed bibliography we refer to the monograph [2]. We can also note that if the field equations do not follow from the unconditional variational principle but admit some additional structures, then a one-way relationship between global symmetries and conservation laws can be possible. For example, a pre-symplectic structure [3], [4] maps symmetries to conserved currents, [5], [6], [7], [8], and a Lagrange anchor [9] defines an inverse mapping from conservation laws to symmetries [10]. In this paper, we construct a generalization of the first Noether theorem, and the converse one, for the case of the conditional extremum problem. We proceed from the usual definitions and assumptions related to the issues of symmetries and conserved currents, known from various textbooks on field theory, see, for example, [11], while for the gauge symmetry we use the standard setup and notation of gauge theories [12].
The article is organised as follows. In the next section, we briefly outline the conditional Lagrange equations that follow from the conditional principle of least action. The conditional extremum of the action is sought in the class of admissible fields that obey differential equations. This Section also addresses the gauge symmetry of the equations that constrain the fields. In Section 3, we consider special infinitesimal global symmetries of the action in the class of the fields subject to differential equations. These conditional symmetries of action represent specializations of gauge transformations of the equations that are imposed to constrain the admissible class of fields. We then show that such conditional global symmetries lead to currents that are conserved at conditional extrema of the action. This fact is a main result of the article. We also prove the converse theorem: Every current, which is conserved at conditional extremals of the action, leads to a conditional symmetry. Throughout this paper, we develop the formalism in its full generality, without restricting the order of derivatives. At the end of Section 3, we provide a brief exposition of the key relations, including an explicit construction of conserved currents for the simplest case where the Lagrangian does not include higher-order derivatives of the fields, and gauge transformations of the constraint equations do not involve higher-order derivatives of the gauge parameters. In Section 4, we illustrate the general formalism of Sections 2-3 by examples. The final section discusses the results and possible areas of application of the proposed formalism, and more broadly, of conditionally Lagrangian theories in general.
2 Field equations for the problem of conditional extremum
@@
$S[\phi(x)]=\int d^{d}xL(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)$ (1)
and impose the differential equations to restrict the admissible fields,
$T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0,\qquad a=1,\ldots,N\,.$ (2)
The problem is to find the minimum of the action (1) among the fields subject to equations (2).
@@
$S_{\lambda}[\phi(x),\lambda(x)]=\int dx\left(L(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)+\lambda^{a}T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)\right)\,.$ (3)
@@
$\frac{\delta S_{\lambda}}{\delta\phi^{i}}=0\,,\qquad\frac{\delta S_{\lambda}}{\delta\lambda^{a}}\equiv T_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0\,$ (4)
@@
$\sum_{k=0}\left(\frac{\partial T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)}{\partial\,\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\phi^{i}}\right)_{\big|T=0}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\delta^{\prime}\phi^{i}=0\,.$ (6)
@@
$\delta^{\prime}S=\int dx\left(\frac{\delta S}{\delta\phi^{i}}\delta^{\prime}\phi^{i}+total\,\,divergence\,\right)=0,$ (7)
and for field configurations admitted by the constraints (2). The above relation does not immediately lead to a system of differential equations, unlike the variational principle for the action with Lagrange multipliers (4), because the variations $\delta^{\prime}\phi^{i}$, being restricted by equations (6), are not arbitrary functions. As a result, the infinitesimal symmetries of the action, even if equations (2) have the same symmetry, do not in general lead to currents that are conserved at the conditional extrema of $S$.
Now let us briefly explain, following [1], how to explicitly construct the field equations that are necessary conditions for the action extremum in the class of fields constrained by differential equations. First, we should find the infinitesimal gauge symmetry transformations for equations (2) which are imposed to select the admissible field configurations. It is assumed that the transformation includes derivatives of the gauge parameters $\epsilon^{\alpha}(x),\alpha=1,\ldots,m$, which are arbitrary smooth functions, up to a certain finite order $K$
$\delta_{\epsilon}\phi^{i}=\hat{R}^{i}_{\alpha}\epsilon^{\alpha}=\sum_{k=0}^{K}\stackrel{{\scriptstyle(k)}}{{R}}{}^{i}_{\alpha}{}^{\mu_{1}\ldots\mu_{k}}(x,\phi,\partial\phi,\partial^{2}\phi\ldots)\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\epsilon^{\alpha}\,$, (8)
where $\stackrel{{\scriptstyle(k)}}{{R}}{}^{i}_{\alpha}{}^{\mu_{1}\ldots\mu_{k}}$ are smooth functions of the fields and their derivatives. The above transformation is supposed to be a symmetry of equations (2), hence
$\delta_{\epsilon}T_{a}{}_{\big|T=0}=0,\qquad\forall\epsilon_{\alpha}(x)\,.$ (9)
@@
$\delta S[\phi(x)]=\int dx\left(\partial_{\mu}J^{\mu}+\frac{\delta S}{\delta\phi^{i}}\delta\phi^{i}\right)\,$, (11)
$J^{\mu}=\sum_{p=1}^{M}p^{\mu\mu_{1}\ldots\mu_{p-1}}_{i}\partial_{\mu_{1}}\ldots\partial_{\mu_{p-1}}\delta\phi^{i}\,$, (12)
$p^{\mu\mu_{1}\ldots\mu_{p-1}}_{i}=\sum_{k=0}^{M-p}(-1)^{k}\partial_{\nu_{1}}\ldots\partial_{\nu_{k}}\frac{\partial L}{\partial(\partial_{\nu_{1}}\ldots\partial_{\nu_{k}}\partial_{\mu}\partial_{\mu_{1}}\partial_{\mu_{p-1}}\phi^{i})}\,$, (13)
@@
$\delta_{\epsilon}S=\int dx\left(\epsilon^{\alpha}\hat{R}^{\dagger i}_{\alpha}\,\frac{\delta S}{\delta\phi^{i}}+\partial_{\mu}I^{\mu}\right)\,$, (14)
@@
$\hat{R}^{\dagger i}_{\alpha}F=\sum_{k=0}^{K}(-1)^{k}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\left(\stackrel{{\scriptstyle(k)}}{{R}}{}^{i\mu_{1}\ldots\mu_{k}}_{\alpha}F\right),$ (15)
@@
$\frac{\delta_{\epsilon}S}{\delta\epsilon^{\alpha}}=\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}=0\,,\hskip 18.49988ptT_{a}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)=0\,.$ (17)
@@
3 Nother’s theorem for the problem of conditional extremum
@@
$\epsilon^{\alpha}_{\omega}(x)=\psi^{\alpha}_{A}(x,\phi,\partial\phi,\partial^{2}\phi,\ldots)\,\omega^{A}\,,\qquad A=1,\dots,s\,$, (18)
@@
$\tilde{\delta}_{\omega}\phi^{i}=\delta_{\epsilon}\phi^{i}_{|\epsilon\,\mapsto\,\epsilon_{\omega}}$ (19)
@@
$\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}\Lambda_{A}^{\mu}(\phi,\partial\phi,\dots)+\hat{W}^{a}_{A}T_{a}\right),$ (20)
@@
$\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}I^{\mu}_{A}+\psi^{\alpha}_{A}\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}\right)\,$, (21)
@@
$\partial_{\mu}\tilde{I}{}^{\mu}_{A}=-\psi^{\alpha}_{A}\hat{R}{}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}+\hat{W}^{a}_{A}T_{a}\,,\qquad\tilde{I}{}^{\mu}_{A}=I^{\mu}_{A}-\Lambda^{\mu}_{A}\,.$ (23)
@@
$\partial_{\mu}{\Phi}{}^{\mu}_{A}=\hat{\Psi}{}^{\alpha}_{A}\hat{R}{}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}+\hat{W}^{a}_{A}T_{a}\,$, (24)
@@
$\hat{\Psi}^{\alpha}_{A}=\sum_{k=0}^{N}\stackrel{{\scriptstyle(k)}}{{\psi}}{}^{\alpha\,\mu_{1}\ldots\mu_{k}}_{A}(\phi,\partial\phi,\partial^{2}\phi,\ldots)\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\,,\hskip 18.49988pt\hat{\Psi}{}^{\alpha}_{A}{}_{\big|T=0}\neq 0,\qquad\forall A\,.$ (25)
@@
$\partial_{\mu}\Phi^{\prime\mu}_{A}=\hat{W}^{a}_{A}T_{a}+\psi^{\alpha}_{A}\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}},$ (26)
@@
$),\displaystyle\Phi^{\prime\mu}_{A}=\Phi^{\mu}_{A}-\sum_{p=0}^{N-1}\bigg(\left(\partial_{\mu_{1}}\ldots\partial_{\mu_{p}}\hat{R}^{\dagger i}_{\alpha}\frac{\delta S}{\delta\phi^{i}}\right)\sum_{k=0}^{N-p-1}\bigg(\partial_{\mu_{p+1}}\ldots\partial_{\mu_{p+k}}\bigg((-1)^{k}\stackrel{{\scriptstyle(k+p+1)}}{{\psi}}{}^{\alpha}_{A}{}^{\mu\mu_{1}\ldots\mu_{p}\mu_{p+1}\ldots\mu_{p+k}}\bigg)\bigg$$\bigg)\,$,
$\psi^{\alpha}_{A}=\sum_{k=0}^{N}(-1)^{k}\partial_{\mu_{1}}\ldots\partial_{\mu_{k}}\stackrel{{\scriptstyle(k)}}{{\psi}}{}^{\alpha}_{A}{}^{\mu_{1}\ldots\mu_{k}}\,$, (28)
@@
$\delta_{\epsilon}\phi^{i}=\hat{R}^{i}_{\alpha}\epsilon^{\alpha},\qquad\epsilon_{\omega}=\psi^{\alpha}_{A}\omega^{A}\,,\qquad\tilde{\delta}_{\omega}\phi^{i}=\delta_{\epsilon_{\omega}}\phi^{i}\,.$ (29)
@@
$\tilde{\delta}_{\omega}L=\omega^{A}\left(\partial_{\mu}\left(I^{\mu}_{A}+\Phi^{\prime\mu}_{A}\right)+\hat{W}{}^{a}_{A}T_{a}\right)\,.$ (30)
@@
The formalism considered in the two previous sections covers the general case where the Lagrangian can include derivatives of fields of any order, and gauge transformations (8) can involve any derivatives of the gauge parameters. The full generality of the formalism may leave the impression that the construction of conditionally Lagrangian equations and the conserved currents is a rather cumbersome procedure. However, Lagrangians without higher-order derivatives are most often studied, and gauge symmetries also most often include only the first derivatives of the parameters. In such a situation, without higher-order derivatives, both the conditional Lagrangian equations and the corresponding conditional symmetries and conserved currents take on a simpler form. Therefore, for the convenience of those readers who might be more interested in searching for conditional conservation laws in first-order models than in the generality of the formalism, we present a summary of the key expressions for this simpler type of theory. In this most common class of Lagrangians and gauge symmetries, the formalism is quite straightforward, and not much more complicated than the corresponding constructions in the unconditional action extremum problem.
Let us assume that gauge transformation (8) of the constraining equations (2) involves at most the first derivatives of the gauge parameters,
$\delta_{\epsilon}\phi^{i}=\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}(\phi,\partial\phi,\partial^{2}\phi,\dots)\epsilon^{\alpha}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}(\phi,\partial\phi,\partial^{2}\phi,\dots)\partial_{\mu}\epsilon^{\alpha}\,,\qquad\delta_{\epsilon}T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)_{\big|T=0}=0\,,\,\,\forall\epsilon^{\alpha}(x).$ (31)
Let us further assume that the Lagrangian does not involve higher derivatives. The conditional extrema of the action are subject to conditional Lagrangian equations (17), which in this case read:
$\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\mu}\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\right)-\partial_{\mu}\left(\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\left(\frac{\partial L}{\partial\phi^{i}}-\partial_{\mu}\frac{\partial L}{\partial\partial_{\mu}\phi^{i}}\right)\right)=0\,,\qquad T_{a}(\phi,\partial\phi,\partial^{2}\phi,\dots)=0\,.$ (32)
@@
$\tilde{\delta}_{\omega}\phi^{i}=\left(\stackrel{{\scriptstyle(0)}}{{R}}{}^{i}_{\alpha}\psi^{\alpha}_{A}+\stackrel{{\scriptstyle(1)}}{{R}}{}^{i}_{\alpha}{}^{\mu}\partial_{\mu}\psi^{\alpha}_{A}\right)\omega^{A}\,.$ (33)
@@
$\tilde{\delta}_{\omega}L_{\big|T=0}=\omega^{A}\partial_{\mu}\Lambda^{\mu}_{A}(\phi,\partial\phi,\dots)\,.$ (34)